If the length of one side of a square is 2h+3, express in terms of h
A. the perimeter of the square
B. The area of the square

Answer by lwsshak3(11628)

If the length of one side of a square is 2h+3, express in terms of h
A. the perimeter of the square = 4*(2h+3)=8h+12
B. The area of the square=(2h+3)^2=4h^2+12h+9